Calculate the boiling point and freezing point of a solution prepared by the solving 68g of glocuse(C6H12O6) in 800g of water. The moral mask of C6H12O6 is 180g/mol​
1
Expert's answer
2020-10-08T13:55:39-0400

The molecular weight of glucose C6​H12O6​=6(12)+12(1)+6(16)=180 g/mol.
The number of moles of glucose =180g/mol
0.625g​=0.00347moles
Mass of water =102.8g=0.1028kgThe molality of glucose m=0.1028kg
0.00347moles​=0.0338mol/kg
The depression in the freezing point ΔTf​=Kf​m=1.87Kkg/mol×0.0338mol/kg=0.0632K
Freezing point of water =273 K
The freezing point of the solution =273+0.0632=273.0632K